# Spectral subspace extraction via incoherent quantum phase estimation

### v1

@@ Abstract
Quantum phase estimation (QPE) is a cornerstone algorithm for extracting Hamiltonian eigenvalues, but its standard form targets individual eigenstates and requires carefully prepared coherent inputs. To overcome these limitations, we adopt an ensemble-based formulation of QPE that estimates the density of states (DOS) of the Hamiltonian generator of the evolution. This approach, which we refer to as DOS-QPE, builds on a prior formulation introduced by one of the authors. In this work, we present DOS-QPE as a circuit primitive, extending it with symmetry-adapted input ensembles and advanced spectrum reconstruction techniques. This variant of QPE enables natural access to thermodynamic properties, symmetry-resolved spectral functions, and features relevant to quantum many-body systems. We demonstrate its performance on fermionic models and nuclear Hamiltonians by casting the spectrum reconstruction problem as a quadratic program solved via compressed sensing. These use cases highlight the potential of DOS-QPE for early fault-tolerant quantum simulations in spectroscopy, electronic structure, and nuclear theory.
@@ I Introduction
Estimating the spectral properties of quantum systems is a fundamental problem with wide-ranging applications in physics, chemistry, and quantum computing. In particular, the density of states (DOS) — the distribution of energy eigenvalues of a quantum Hamiltonian — encodes crucial information about both the thermodynamic behavior [1] and the dynamical response of a system [2]. Beyond its traditional role in condensed matter physics and statistical mechanics, the DOS has recently been used in hybrid classical-quantum frameworks for quantum topological data analysis [3] and to address the frustration index problem in signed graphs [4].
Quantum phase estimation (QPE) remains the gold standard for extracting eigenvalue information in the form of quantum phases [5, 6, 7, 8]. It is typically applied to eigenstates of unitary evolution operators or their generators, and its success depends on the overlap between the prepared trial state and the true eigenstate of the Hamiltonian. However, this requirement can be prohibitive in practice. Preparing trial states with significant overlap may demand deep, problem-specific quantum circuits, especially for strongly-correlated systems [9, 10, 11]. In addition, real quantum hardware inevitably interacts with its surrounding environment, complicating further the preparation of single pure states. This raises the natural question of how compatible QPE is with mixed states, thermal states, and noisy, non-pure ensembles [12]. While purification techniques [13, 14] are well-established in quantum information, the use of QPE on purified states has not yet been systematically studied as a circuit primitive. Notable exceptions include the use of garbage states in quantum topological data analysis [15] and recent work leveraging QPE on maximally mixed states to compute Hamiltonian moments in the context of signed graphs [4].
While a promising alternative approach for estimating the DOS using Chebyshev moments has been recently proposed [16], this work focuses on developing a framework centered around QPE. This not only provides access to a rich landscape of algorithmic variants, but also positions our method for reuse as a subroutine in a range of QPE-based quantum algorithms—including Shor’s algorithm [17], quantum simulation in quantum chemistry [18], and the HHL algorithm for solving linear systems [19].
In this work, we generalize and refine a QPE-based method for DOS estimation (DOS-QPE) [4], presenting it as a circuit primitive. We showcase some of its key applications, ranging from the Fermi-Hubbard model to molecular systems and nuclear physics Hamiltonians. We begin with a review of standard QPE techniques and the modifications required to apply them to purified states. We then discuss suitable mixed state preparations and rescaling of the Hamiltonian. We propose a method to resolve the problem of integer degeneracies and, more generally, tackle spectrum reconstruction using the tools from quadratic programming [20] and compressed sensing [21, 22]. Finally, we present three case studies of the protocol: the Fermi-Hubbard Hamiltonian, a molecular electronic structure problem, and a strongly correlated nuclear physics Hamiltonian.
@@ II DOS-QPE
Standard QPE is the fault-tolerant benchmark for the estimation of phases (eigenvalues) in the case of a pure, coherent eigenstate of the unitary iiiThroughout this paper, we refer to coherent and incoherent in the sense of resource theories [23, 24]. Specifically, we refer to states as coherent when their density matrix contains non-zero off-diagonal elements (coherences), and as incoherent when these off-diagonal terms vanish.. Its successful execution requires the prepared probing state to have substantial overlap with the target eigenstate whose phase we aim to measure. Although several techniques have been proposed to address the state preparation problem, this step often incurs significant overhead and is sometimes the most challenging part of the protocol.
In chemistry, trial eigenstate preparation remains a hot topic [25]; for weakly correlated systems, the qubit-mapped Hartree-Fock state may be good enough [26], but for more strongly correlated systems, elaborate state preparation protocols are required. These usually take the form of a variational quantum algorithm [27], which represent difficult optimization problems [28, 29] and suffer prohibitive sampling costs [30]. Otherwise, approximate classical eigenstates can be loaded onto a qubit register by an appropriate mapping such as CVO-QRAM [31]. Still, even this approach does not scale favorably with system size, not to mention the additional classical cost of finding a high-quality approximation to the target eigenstate, using, for example, one of the selected CI methods [32, 33, 34]. In addition, we might be interested in more general properties of the Hermitian operator generator of the unitary evolution. For example, accessing information encoded in higher-energy states may require reconstructing the entire spectrum of the target Hamiltonian. This is the case we will consider in the following.
We are interested in the reconstruction of the so-called eigenvalue function [35] or density of states of the Hamiltonian operator H{H}. The density of states (DOS) plays a central role in extracting spectral properties of an operator. It is typically derived from the Fourier transform of the operator’s unitary time evolution and is essential for the reconstruction of the microcanonical ensemble of the system [36]. The DOS technique is a flexible approach for gathering information about both low- and high-frequency components.
@@
To implement this approach into standard QPE (see App. A for the derivation), we need to add an additional ancillary register that we will refer to as the purification register. This register matches the state register dimension, but it will be traced out at the end of the protocol. By means of this operation, we are effectively evaluating the trace over the system $P$ in Eq. (2). This will allow us to translate QPE from a coherent to an incoherent approach, extending the routine to support further practical applications. Note that purification is not the only way to transition between coherent and incoherent QPE. In case of interest in statistical properties of the systems, any thermal state preparation routine [37, 38, 39] will suffice as long as the output state can be matched to the state register.
@@
II.1 Maximally mixed state
The most standard example of purification is that of the maximally mixed state [13]. This purification is easily achieved via the maximally entangled state
$|\Phi\rangle_{SP}=\frac{1}{\sqrt{2^{n}}}\sum_{k=1}^{2^{n}}|k\rangle_{S}\otimes|k\rangle_{P},$ (6)
where $\{|k\rangle_{P}\}$ is a basis for the auxiliary system $\mathcal{H}_{P}\cong\mathcal{H}_{S}$. Here and in the following, $S$ and $P$ indicate the state and the purification registers. Tracing out the system $P$, we have
[…]
that is, an ensemble of states covering the entire Hilbert space. Using this ensemble with flat distribution, we can probe all of the Hamiltonian’s eigenvalues (phases) with equal probability and estimate the Hamiltonian’s spectral resolution. The power and potential of using maximally mixed states in QPE have been shown, under the umbrella of garbage state preparation and kernel extraction, in the original proposal of quantum topological data analysis by Lloyd et al. [15]. In a similar way, but with the addition of the circuit primitive, DOS-QPE has been shown capable of generating features required for learning the frustration index [4].
II.2 Dicke states
Given some $M$-dimensional set of discrete fermionic modes corresponding to single-particle quantum states, we can write the resulting $2^{M}$-dimensional Fock space as a direct sum of subspaces $\mathcal{F}_{k}$,
$\mathcal{F}=\bigoplus_{k}\mathcal{F}_{k}=\mathcal{F}_{0}\oplus\mathcal{F}_{1}\oplus\cdots\oplus\mathcal{F}_{M}.$ (8)
@@ II.2 Dicke states
Each term in the decomposition shown in Eq. (8) corresponds to the subspace spanned by the set of occupation number vectors with a fixed particle number, $k$, of which there are $\binom{M}{k}$. It is typically the case for fermionic Hamiltonian simulations that the eigenvalues and eigenvectors of interest all carry the same particle number. As such, using the maximally mixed state as the probe for DOS-QPE may lead to a spectrum that is too crowded for certain applications or lead to a needlessly high sampling cost.
A simple way to enforce particle-number symmetry in the spectrum obtained from DOS-QPE is to choose a fermion-qubit encoding which maps fermionic states to qubit states conserving the Hamming weight of the original fermionic state after transformation, i.e., the same number of “1”s (hot-entries) in both fermionic and qubit states. The two projections of a single qubit represent a state that is either empty or occupied by a particle, in this case. The Jordan-Wigner encoding [40] satisfies these requirements,
$|x_{0},x_{1},\ldots,x_{M}\rangle\rightarrow|q_{0},q_{1},\ldots,q_{n}\rangle,$ (10)
where the number of qubits, $n$, is the same as the number of fermionic modes, $M$ and $x_{n}=q_{n}\in\{0,1\}$. To restrict the eigenvalues to the correct subspace, then, becomes an issue of preparing, in an even superposition, all computational basis states with Hamming weight $k$ on the state register before purification. Such states are known as Dicke states, $|D^{n}_{k}\rangle$,
@@
where $n$ is the number of qubits, and $k$ is the Hamming weight, which according to the literature can be prepared deterministically with circuit depths of $\mathcal{O}(k\log\frac{n}{k})$, $\mathcal{O}(k\sqrt{\frac{n}{k}})$ and $\mathcal{O}(n)$ on devices with all-to-all, grid and linear nearest neighbour connectivities, respectively [41].
Perhaps the most natural application of DOS-QPE with a Dicke state probe is the extraction of eigenspectra of many-body Hamiltonians. In standard QPE and many of its extensions, a trial state with a large overlap with the target eigenstate must be found either classically or with some other quantum algorithm, such as VQE. Preparing such trial states with a variational quantum algorithm can be prohibitively expensive in terms of sampling costs. Additional difficulties during optimization arise from barren plateaus, and it is not possible to know in advance whether the prepared state has a large enough overlap with the target state, except for some heuristic cases. Preparing a state classically and compiling it into a state-preparation circuit is often more practical, though it still poses challenges, particularly in efficiently encoding the classical state onto the state register.
In contrast, Dicke-probed DOS-QPE does not require preparation of fermionic eigenstates, preserves particle-number symmetry under the Jordan-Wigner mapping, and experiments need to be set up only once to recover the full eigenspectrum. The classical cost incurred before the execution of the algorithm in the simplest case is just the cost of a single Hartree-Fock calculation, which scales polynomially in system size with even naive implementations, plus the cost of constructing the circuit.
@@ III Spectrum reconstruction
Transitioning from estimating a single phase using QPE to reconstructing the full spectrum with DOS-QPE requires a more refined post-processing step. This additional cost is justified by the richer information retrieved from he spectrum. DOS-QPE effectively samples from the normalized density of states,
@@ III.1 The degeneracy integer problem
With access to the readout of the time-frequency register of the DOS-QPE circuit, we aim to estimate the spectral resolution, ${H}=\sum_{k}\theta_{k}|k\rangle\langle k|$. To do so, we face the problem of generating the best spectral estimate out of the readout probability distribution $P$. We refer to this estimation task as the degeneracy integer problem, falling under the umbrella of spectral density estimation [42].
@@ III.2 Compressed sensing and optimization
We formulate the problem on a grid of candidate phases $\{\theta^{\prime}_{j}\}\subset[0,1)$, finer than the hardware-limited resolution $2^{-m}$. We obtain an underdetermined system where the number of potential spectral components exceeds the number of measured bins. To select a physically meaningful sparse solution, we employ a quadratic program [43] with an $\ell_{1}$ regularization penalty (basis-pursuit, LASSO) [44], which encourages sparsity by penalizing the absolute sum of the degeneracy coefficients. This results in a practical realization of compressed sensing [21, 22], a framework for recovering sparse signals from limited or noisy measurements. The finer grid provides an overcomplete dictionary of possible phases, while the $\ell_{1}$ penalty enforces that only a small subset is selected. The resulting optimization routine enables a super-sampled spectral reconstruction despite a coarser hardware discretization [45].
@@
The problem naturally maps onto a mixed-integer quadratic program [46, 47] with convex, non-linear objective function $L_{2,\text{MI}}=\min_{\mathbf{w}\in\mathbb{N}_{+}}|A\mathbf{w}-\mathbf{P}|_{2}^{2}$ with $\sum_{j=1}^{G}w_{j}=2^{n}$. Quadratic programming is NP-complete [48] in general, with the variant of mixed-integer programming (MIP) enforcing $w_{j}\in\mathbb{Z}+$ and $\sum_{j}w_{j}=\sum_{i}d_{i}=2^{n}$ being NP-hard [49, 50, 51]. For this reason, we relax the problem from integer degeneracies to real and add a further final step of clustering of phases and integer approximation on the solution. This makes the problem a quadratic program with a convex objective function, which, together with linear programs, is in P [52, 50, 51]. The final objective function we optimize against is
@@
where we added the $\ell_{1}$ regularization penalty controlled by $\lambda>0$. The squared-$L_{2}$ term ensures data fidelity, and the $\ell_{1}$ term promotes sparse support in the candidate-phase weights. This convex QP can be solved efficiently at moderate $G$ using off-the-shelf solvers. In our simulations, we do this using the CLARABEL solver [53] and CVXPY [54, 55].
@@
To quantify the reconstruction error, we use the 1-Wasserstein distance [56] $W_{1}(\mu,\nu)$ between discrete spectral distributions $\mu=\sum_{i}d_{i}\,\delta_{\theta_{i}}$ and $\nu=\sum_{j}\hat{d}_{j}\,\delta_{\hat{\theta}_{j}}$, where $\delta_{\theta}$ denotes the Dirac measure at $\theta$. This metric is defined as
@@
where $\Gamma(\mu,\nu)$ is the set of all couplings between $\mu$ and $\nu$ that preserve total weight. Intuitively, this metric quantifies the dissimilarity between the two distributions by measuring the minimum amount of ”mass” times ”travel” that needs to be moved to shape one into the other [57]. In the following, we compute $W_{1}$ between the sampled spectrum and the exact one, as well as between the optimized solution and the exact spectrum. This provides a principled way to evaluate the fidelity of spectral estimation, taking into account both eigenphase positions and degeneracies.

### v2

# Purified phase estimation samples spectra efficiently

@@ Abstract
Quantum phase estimation (QPE) is a cornerstone algorithm for extracting Hamiltonian eigenvalues, but its standard, eigenstate-centric form relies on carefully prepared coherent inputs that are costly or impractical for many strongly correlated systems. We overcome this bottleneck via DOS-QPE, an incoherent, purification-based variant of QPE that works directly with mixed-state probes and estimates the density of states (DOS) of the Hamiltonian. By adding a purification register and simple entangling layers, we turn standard QPE into an ensemble-based DOS sampler without modifying the core phase-estimation block. Conceptually, this purification closely aligns with the recent random purification channel framework from quantum learning theory, but instantiated here as a concrete phase-estimation circuit. We further equip DOS-QPE with symmetry-adapted input ensembles and a compressed-sensing reconstruction pipeline, and demonstrate on fermionic and nuclear Hamiltonians that a single experimental setup can recover rich spectral information relevant to thermodynamics, spectroscopy, and many-body structure.
@@ I Introduction
Quantum phase estimation (QPE) remains the gold standard for extracting eigenvalue information in the form of quantum phases [kitaev1995quantum, Abrams_1997, Abrams_1999, dobvsivcek2007arbitrary]. In its textbook formulation, QPE is an eigenstate-centric protocol: one supplies a carefully prepared pure state with large overlap on a target eigenvector, and the circuit returns an estimate of the associated phase. For strongly correlated or structurally complex Hamiltonians, however, preparing such trial states may require deep, problem-specific circuits or demanding variational algorithms [plesch2011quantum, sun2023asymptotically, zhang2022quantum]. On real devices, noise and decoherence further erode the assumption that a single, well-isolated eigenstate can be reliably prepared. This motivates a different question: can we use the machinery of phase estimation to interrogate entire spectral subspaces via mixed or ensemble states instead of finely tuned eigenstates?
Purification offers a natural route to such an “incoherent” formulation of QPE. Any mixed state can be realized as the reduced state of a larger pure state on a system plus auxiliary register [Nielsen_2012, Chiribella_2010]. Conceptually, nothing prevents us from feeding this purification into a QPE circuit and tracing out the purifying system at the end. Despite the ubiquity of purifications in quantum information, this perspective has not been developed as a general circuit primitive for phase estimation. Existing uses are scattered across specific settings, such as garbage states in quantum topological data analysis [Lloyd2016] or maximally mixed probes for Hamiltonian moments in signed graphs [scali2025signed], rather than forming a unified framework for ensemble-based spectral extraction.
In this work we take purification as the organizing principle for QPE and build a density-of-states variant, DOS-QPE, around it. By adding a dedicated purification register and preparing controlled entanglement between the system and its purifier, we turn standard QPE into an incoherent primitive that samples the density of states of the Hamiltonian generator. The QPE block itself remains essentially unchanged; what we redesign is the input. Instead of a single pure eigenstate, we supply mixed or symmetry-adapted ensembles implemented via simple purification circuits. This shift allows us to access thermodynamic and spectral information directly from ensembles, and to probe entire invariant subspaces without ever preparing individual eigenstates.
While developing this purified-QPE framework, we became aware of concurrent progress in quantum learning theory that points in a similar direction. The line of work by Tang et al. [tang] and Pelecanos et al. [pelecanos] shows that mixed-state learning and tomography can be efficiently reduced to pure-state problems via a random purification channel (RPC), an instance of the “acorn trick” converting multiple copies of a mixed state into random purifications suitable for pure-state algorithms. Girardi et al. [girardi] further provide an explicit and simple construction of the RPC. These results offer a natural theoretical backdrop: they identify purification not as a technical afterthought, but as a fundamental algorithmic resource. In hindsight, DOS-QPE can be viewed as the phase-estimation analogue of this sample-efficient paradigm, leveraging purification to turn incoherent ensembles into usable input for a standard, well-studied quantum subroutine.
We deliberately base our DOS estimation framework on QPE rather than on alternative approaches such as Chebyshev-moment methods [Summer_2024]. QPE sits at the core of many quantum algorithms, and recasting it in an incoherent, purification-based form preserves compatibility with this broader ecosystem. Our primitive can in principle be used as a drop-in component in QPE-based routines, including Shor’s algorithm [Shor], quantum simulation for quantum chemistry [Aspuru_Guzik_2005], and the HHL algorithm for solving linear systems [Harrow_2009], enabling ensemble-based versions of otherwise eigenstate-driven procedures.
In this work, we generalize and refine a QPE-based method for DOS estimation (DOS-QPE) [scali2025signed], and elevate it to a reusable circuit primitive. We (i) formalize DOS-QPE in a purified setting, making explicit how a purification register and simple entangling circuits turn standard QPE into an incoherent DOS sampler; (ii) introduce mixed-state probes tailored to specific symmetry sectors, including maximally mixed ensembles and particle-number–resolved probes constructed from Dicke states; and (iii) develop a convex-optimization–based pipeline for spectrum reconstruction using quadratic programming and compressed sensing [Cornu_jols_2018, Donoho_2006, Cand_s_2006]. We then demonstrate the reach of this approach on three representative classes of Hamiltonians: a Fermi–Hubbard model, an electronic-structure Hamiltonian, and a strongly correlated nuclear Hamiltonian. Together, these examples showcase DOS-QPE as a practical and versatile primitive for spectral subspace extraction, highlighting its potential as a core tool for early fault-tolerant quantum simulations.
@@ II DOS-QPE
Standard QPE is the fault-tolerant benchmark for the estimation of phases (eigenvalues) in the case of a pure, coherent eigenstate of the unitary iiiThroughout this paper, we refer to coherent and incoherent in the sense of resource theories [aberg, Baumgratz_2014]. Specifically, we refer to states as coherent when their density matrix contains non-zero off-diagonal elements (coherences), and as incoherent when these off-diagonal terms vanish.. Its successful execution requires the prepared probing state to have substantial overlap with the target eigenstate whose phase we aim to measure. Although several techniques have been proposed to address the state preparation problem, this step often incurs significant overhead and is sometimes the most challenging part of the protocol.
In chemistry, trial eigenstate preparation remains a hot topic [lee2023evaluating]; for weakly correlated systems, the qubit-mapped Hartree-Fock state may be good enough [tubman2018postponing], but for more strongly correlated systems, elaborate state preparation protocols are required. These usually take the form of a variational quantum algorithm [peruzzo2014variational], which represent difficult optimization problems [larocca2024review, astrakhantsev2023phenomenological] and suffer prohibitive sampling costs [gonthier2022measurements]. Otherwise, approximate classical eigenstates can be loaded onto a qubit register by an appropriate mapping such as CVO-QRAM [deveras2022double]. Still, even this approach does not scale favorably with system size, not to mention the additional classical cost of finding a high-quality approximation to the target eigenstate, using, for example, one of the selected CI methods [huron1973iterative, tubman2016deterministic, sharma2017semistochastic]. In addition, we might be interested in more general properties of the Hermitian operator generator of the unitary evolution. For example, accessing information encoded in higher-energy states may require reconstructing the entire spectrum of the target Hamiltonian. This is the case we will consider in the following.
We are interested in the reconstruction of the so-called eigenvalue function [osborne2006renormalisationgroup] or density of states of the Hamiltonian operator H{H}. The density of states (DOS) plays a central role in extracting spectral properties of an operator. It is typically derived from the Fourier transform of the operator’s unitary time evolution and is essential for the reconstruction of the microcanonical ensemble of the system [kardar]. The DOS technique is a flexible approach for gathering information about both low- and high-frequency components.
@@
To implement this approach into standard QPE (see App. A for the derivation), we need to add an additional ancillary register that we will refer to as the purification register. This register matches the state register dimension, but it will be traced out at the end of the protocol. By means of this operation, we are effectively evaluating the trace over the system $P$ in Eq. (2). This will allow us to translate QPE from a coherent to an incoherent approach, extending the routine to support further practical applications. Note that purification is not the only way to transition between coherent and incoherent QPE. In case of interest in statistical properties of the systems, any thermal state preparation routine [gilyen, Ding_2024, Ding_2024_2] will suffice as long as the output state can be matched to the state register.
@@
II.1 Mixed-state probes
In DOS-QPE, the choice of input ensemble determines which part of the spectrum is probed and with what weights. Rather than relying on a single coherent eigenstate, we work with mixed states
$|,\rho_{S}=\sum_{j}p_{j}\outerproduct{\psi_{j}}{\psi_{j}}$, (6)
@@ II.1 Mixed-state probes
Operationally, this corresponds to adding an extra purification register, preparing an entangled state between $S$ and $P$ with a simple circuit primitive, and then discarding $P$ after QPE. In this work, we focus on two practically relevant classes of mixed states: (i) the maximally mixed state, which leads to a flat sampling over the entire Hilbert space, and (ii) particle-number–resolved mixed states obtained from Dicke-state probes, which restrict DOS-QPE to a fixed Hamming-weight sector.
These ensembles can be realized in two complementary ways. First, we introduce an explicit circuit construction based on an additional purification register and a layer of CNOT gates in cascade that prepares a maximally entangled state between $S$ and $P$, which is then traced out. Second, one may employ the recent Haar-random purification methods developed in the context of quantum learning theory by Tang et al. [tang], Pelecanos et al. [pelecanos], and Girardi et al. [girardi]. There, the random purification channel (RPC) efficiently converts multiple copies of a mixed state into Haar-random purifications that can be processed by pure-state algorithms. In our setting, such Haar-random purifications provide an alternative, theoretically grounded route to preparing mixed-state probes compatible with DOS-QPE without tailoring a problem-specific purification circuit.
II.1.1 Maximally mixed state via Hadamard layer
The most standard example of a mixed-state probe is the maximally mixed state on the $n$-qubit state register,
$\rho_{S}=\frac{\mathbb{I}}{2^{n}},$ (8)
which assigns equal weight to all computational basis states and therefore to all eigenstates of the Hamiltonian under a generic encoding. A convenient purification of $\rho_{S}$ is the maximally entangled state
@@ II.1.1 Maximally mixed state via Hadamard layer
[…]
that is, the maximally mixed state on $\mathcal{H}_{S}$.
[…]
After running QPE with controlled-$U_{H}^{k}$ acting only on $S$, we trace out (or equivalently, discard) the purification register $P$, thereby realizing DOS-QPE with a maximally mixed probe. Using this ensemble with flat distribution, we can probe all of the Hamiltonian’s eigenvalues with equal probability and estimate the Hamiltonian’s spectral resolution. The power and potential of maximally mixed states in QPE have been demonstrated under the umbrella of garbage state preparation and kernel extraction in quantum topological data analysis [Lloyd2016], and, with the addition of the DOS-QPE circuit primitive, in generating features for learning the frustration index in signed graphs [scali2025signed].
II.1.2 Particle-number–resolved mixed states via Dicke probes
In many fermionic simulations, the eigenstates of interest lie within a fixed-particle-number sector. To exploit this structure, we consider mixed-state probes supported only on the subspace of fixed Hamming weight $k$. Let $\mathcal{F}$ be the $2^{M}$-dimensional Fock space associated with $M$ discrete fermionic modes, which decomposes as
$\mathcal{F}=\bigoplus_{k=0}^{M}\mathcal{F}_{k},$ (12)
[…]
with Hamming weight $|\bm{x}|=k$.
It is often the case that the physically relevant spectrum is contained in a fixed $\mathcal{F}_{k}$. If we were to use the maximally mixed state on the full Hilbert space as the DOS-QPE probe, the resulting spectrum may be unnecessarily crowded and incur a large sampling overhead, as many samples would fall outside the sector of interest. A simple way to enforce particle-number symmetry in DOS-QPE is to choose a fermion–qubit encoding that preserves Hamming weight, such as the Jordan–Wigner mapping [JordanWigner],
[…]
with $n=M$ and $x_{i}=q_{i}\in\{0,1\}$. Under this encoding, the two projections of each qubit represent empty or occupied fermionic modes.
To restrict the probe to the correct subspace, we prepare an even superposition over all computational basis states of Hamming weight $k$ on the state register before purification. These are the Dicke states,
[…]
Perhaps the most natural application of DOS-QPE with a Dicke-state probe is the extraction of eigenspectra of many-body Hamiltonians at fixed particle number. In standard QPE and many of its extensions, one must prepare a trial state with large overlap with a particular eigenstate, often through variational algorithms such as VQE, which are costly to optimize and may suffer from barren plateaus. In contrast, Dicke-probed DOS-QPE does not require preparation of individual fermionic eigenstates, preserves particle-number symmetry under Jordan–Wigner, and allows a single experimental configuration to recover the full eigenspectrum within the chosen particle-number sector. The dominant classical cost in the simplest workflow is that of a single Hartree–Fock calculation, which scales polynomially with system size even for naive implementations, plus the cost of constructing the corresponding quantum circuit.
@@ III Spectrum reconstruction
Transitioning from estimating a single phase using QPE to reconstructing the full spectrum with DOS-QPE requires a more refined post-processing step. This additional cost is justified by the richer information retrieved from the spectrum. DOS-QPE effectively samples from the normalized density of states,
@@ III.1 The degeneracy integer problem
With access to the readout of the time-frequency register of the DOS-QPE circuit, we aim to estimate the spectral resolution, ${H}=\sum_{k}\theta_{k}|k\rangle\langle k|$. To do so, we face the problem of generating the best spectral estimate out of the readout probability distribution $P$. We refer to this estimation task as the degeneracy integer problem, falling under the umbrella of spectral density estimation [stoica2005spectral].
@@ III.2 Compressed sensing and optimization
We formulate the problem on a grid of candidate phases $\{\theta^{\prime}_{j}\}\subset[0,1)$, finer than the hardware-limited resolution $2^{-m}$. We obtain an underdetermined system where the number of potential spectral components exceeds the number of measured bins. To select a physically meaningful sparse solution, we employ a quadratic program [Boyd_2004] with an $\ell_{1}$ regularization penalty (basis-pursuit, LASSO) [Hastie_2015], which encourages sparsity by penalizing the absolute sum of the degeneracy coefficients. This results in a practical realization of compressed sensing [Donoho_2006, Cand_s_2006], a framework for recovering sparse signals from limited or noisy measurements. The finer grid provides an overcomplete dictionary of possible phases, while the $\ell_{1}$ penalty enforces that only a small subset is selected. The resulting optimization routine enables a super-sampled spectral reconstruction despite a coarser hardware discretization [Bertsimas_2024].
@@
The problem naturally maps onto a mixed-integer quadratic program [Nesterov_2004, Conforti_2014] with convex, non-linear objective function $L_{2,\text{MI}}=\min_{\mathbf{w}\in\mathbb{N}_{+}}|A\mathbf{w}-\mathbf{P}|_{2}^{2}$ with $\sum_{j=1}^{G}w_{j}=2^{n}$. Quadratic programming is NP-complete [Vavasis_1990] in general, with the variant of mixed-integer programming (MIP) enforcing $w_{j}\in\mathbb{Z}+$ and $\sum_{j}w_{j}=\sum_{i}d_{i}=2^{n}$ being NP-hard [Pia_2016, Renegar_1992, Liberti_2019]. For this reason, we relax the problem from integer degeneracies to real and add a further final step of clustering of phases and integer approximation on the solution. This makes the problem a quadratic program with a convex objective function, which, together with linear programs, is in P [Kozlov_1980, Renegar_1992, Liberti_2019]. The final objective function we optimize against is
@@
where we added the $\ell_{1}$ regularization penalty controlled by $\lambda>0$. The squared-$L_{2}$ term ensures data fidelity, and the $\ell_{1}$ term promotes sparse support in the candidate-phase weights. This convex QP can be solved efficiently at moderate $G$ using off-the-shelf solvers. In our simulations, we do this using the CLARABEL solver [clarabel] and CVXPY [cvxpy, cvxpy2].
@@
To quantify the reconstruction error, we use the 1-Wasserstein distance [Kantorovich_1960] $W_{1}(\mu,\nu)$ between discrete spectral distributions $\mu=\sum_{i}d_{i}\,\delta_{\theta_{i}}$ and $\nu=\sum_{j}\hat{d}_{j}\,\delta_{\hat{\theta}_{j}}$, where $\delta_{\theta}$ denotes the Dirac measure at $\theta$. This metric is defined as
@@
where $\Gamma(\mu,\nu)$ is the set of all couplings between $\mu$ and $\nu$ that preserve total weight. Intuitively, this metric quantifies the dissimilarity between the two distributions by measuring the minimum amount of “mass” times “travel” that needs to be moved to shape one into the other [Villani_2009]. In the following, we compute $W_{1}$ between the sampled spectrum and the exact one, as well as between the optimized solution and the exact spectrum. This provides a principled way to evaluate the fidelity of spectral estimation, taking into account both eigenphase positions and degeneracies.